OpenStudy (anonymous):

2x - 2y + 5z - 2x - y + 3z x - 3y + 8z 2x - 2y + 8z 3y + 8z -3y + 8z

OpenStudy (anonymous):

-3y+8z

OpenStudy (anonymous):

switch like terms next to each other (2x - 2y + 5z - 2x - y + 3z) change to ( 2x - 2x + y - 2y - 5z + 3z) next the y thats all alone add one in front of it.. thats just saying that theres only one y add/subtract like terms 2x - 2x = 0, 1y - 2y = 1y, 5z + 3z = 8z so you'll have ( 0 + 1y - 8z) leave the zero and youll answer : 1y - 8z

OpenStudy (anonymous):

It's -2y -y You add

OpenStudy (anonymous):

-3y + 8z

OpenStudy (anonymous):

Thanks You Guys !